# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

### v2

@@ 4.3 Baselines
In the Mind2Web dataset, previous work has not treated it as how-to questions. State-of-the-art methods like SYNASE (Zheng et al., 2023) and MINDACT (Deng et al., 2023a) either use In-Context Learning (ICL), providing few-shot demonstrations, or Supervised Learning (SL) to finetune a model on the training set. To ensure a fair comparison with our LLM endpoints [^7] We use GPT-3.5 and GPT-4 with version 1106-preview., we re-implement the MINDACT experiments with the same demonstrations and included chat history as extra context. In our paper, we treat the Mind2Web dataset as dynamic how-to questions and use our RAG system to solve these tasks. For comparison, we use doc-based (providing the entire document) and chunk-based data organization paradigms as baselines against Thread. This RAG system with different data organization paradigms also serves as baselines for the WikiHow and IcM datasets. More details about experiments can be found in Appendix LABEL:appendix_expD.
4.4 Evaluation Metrics
For the Mind2Web dataset, each task is treated as a multi-choice question. We adapt the evaluation metrics used in (Deng et al., 2023a), which include: Element Accuracy (Ele. Acc) to compare the chosen HTML element with all provided elements; Operation F1 (Op. F1) to calculate the token-level F1 score for predicted operations such as “CLICK”, “TYPE IN”, etc.; Step Success Rate (Step SR), where a step is successful if both the selected element and predicted operation are accurate; and Success Rate (SR), where a task is successful only if all steps are successful.
For the WikiHow dataset, which contains ground truth steps, we leverage LLMs to extract “Action Items” from each ground truth step and generated step, and we use the following metrics: Precision ($P=\frac{\#matched\_items}{\#total\_generated\_items}$); Recall ($R=\frac{\#matched\_items}{\#total\_groundtruth\_items}$); F1; and Success Rate (SR) to assess if the generated steps can successfully complete the task, using LLMs to evaluate (Appendix C.3 shows the evaluation prompt).
For the IcM dataset, which involves task execution, we perform evaluations with OCEs (refer to §4.1) using five metrics: Success Rate (SR) indicating the percentage of incidents mitigated automatically by the system without human intervention; Step Success Rate (Step SR) representing the percentage of successful steps out of all task steps; Pre-Failure Step Success Rate (P.F. Step SR) representing the percentage of successful steps before the first failure; Human Intervention (HI) measuring the percentage of steps requiring human intervention; and Average Turns (Turns) to measure the average interaction turns between OCEs and the system during incident mitigation.
5 Experimental Results
5.1 Main Results
Web Navigation. The overall performance on Mind2Web is shown in Table 2, where we compare our method with baselines and both doc-based and chunk-based RAG methods. We observe that providing informative documents (regardless of the data organization paradigm) within RAG methods significantly improves performance. RAG methods outperform ICL and SL methods. By incorporating external documents, the doc-based RAG method achieves performance comparable to the best results of MINDACT. Notably, our Thread paradigm is the best among RAG methods, showing improvements of 4.06% in Ele. Acc, 3.49% in Step SR and 3.57% in SR. Note that MINDACT-SL gets the highest Op. F1 due to label distribution imbalance [^8] If the model predicts all operations as ”CLICK,” the Op. F1 would reach 79.90%., leading the model to favor generating the most frequent operations.
@@ 5.1 Main Results
Table 2: Experiment results on Mind2Web. ‘*’ represents taking results from the original paper.
@@
Paradigm Category SR Step SR P.F. Step SR HI Turns
Simple 40.51 60.90 60.90 30.10 3.14
Chunk Hard 28.95 53.16 43.05 46.84 6.84
Simple 43.86 63.90 63.90 36.09 2.98
Doc Hard 31.58 57.89 42.11 42.11 6.53
Simple 77.19 88.72 84.21 11.28 2.56
Thread Hard 52.63 84.21 68.95 15.79 5.74
Table 3: Experiment results on Incident Mitigation. We divide incidents into two groups: simple or hard, and compare the performance separately.
Incident Mitigation. Table 3 illustrates the advantages of Thread over other paradigms when addressing complex dynamic how-to questions. Both chunk-based and doc-based RAG methods exhibit significant limitations in incident mitigation, as their SR and P.F. Step SR are much lower due to their inability to connect subsequent steps based on the current step. In contrast, our Thread achieves the best score across all metrics, particularly achieving a significant increase in SR from 21.02% to 33.33%. More importantly, the highest P.F. Step SR highlights Thread’s ability to dynamically connect to subsequent steps based on user feedback, thereby reducing the need for human intervention. Consequently, Thread not only achieves the highest Step SR but also requires the fewest interaction turns to effectively mitigate both simple and complex incidents.
@@
Table 4: Experiment results on WikiHow with different interaction manners and paradigms.
Wikipedia Instructions. Table 4 presents the performance of various data organization paradigms on WikiHow tasks, experimenting with both single-turn and multi-turn interactions. In the single-turn setting, where the RAG system needs to generate the entire plan in a single interaction, our Thread outperforms the doc-based method with 5.15%. This indicates that Thread delivers information with less redundancy while achieving a higher SR score. Additionally, when comparing Thread to the chunk-based method, we observe a significantly lower success rate (SR) of 19.59%. This result highlights the challenge of retrieving all relevant chunks without considering the logic flow inside documents.
In the multi-turn setting, the RAG system performs iterative retrieval, showing superior performance compared to single-turn. This indicates the advantage of a step-by-step approach in handling how-to questions. We can see that the SR improves significantly from 58.76% to 68.04% in the doc-based paradigm and from 63.91% to 72.16% in Thread. Similar to the single-turn setting, the chunk-based method splits documents while disrupting internal logic, resulting in a low SR of 20.62%. Our Thread excels across all metrics, achieving the highest SR of 72.16%, highlighting its effectiveness in maintaining and modeling the connections between steps.
@@ 5.2 Ablation on RAG System Settings
This section presents an ablation study on the key settings of our RAG system. While retriever and generator variants have been explored in other research (Gao et al., 2023), we use text-embedding-ada-002 (OpenAI, 2022) for the retriever and GPT-4 as the LLM-based generator. The ablation study is performed on Mind2Web dataset.
Multi-turn Interaction. As shown in Table 4, the multi-turn setting is better than the single-turn setting in answering how-to questions. We use multi-turn setting for the other scenarios in this paper.
Chat History. The chat history helps the system to check previous actions and results, helping the system make better decisions. Without chat history, the performance drops as shown in Table 5. We include chat history for all RAG-based methods in our experiments.
@@
Retrieval Units Selector. As aforementioned in §3.4 and Figure 3, the Selector picks the most relevant retrieval units from the top-K retrieved units. Table 5 shows the ablation of the retrieval unit selector. Comparing chunk selection to LU selection, the performance of Ele. Acc drops by 4.29% and Op.F1 by 3.38%. Conversely, the selector improves all metrics with applied to Thread. Unlike LU selection, which filters out irrelevant LUs based on prerequisites, chunk selection disregards inter-chunk connections and may exclude relevant chunks. Therefore, we activate the selector only for our Thread.
5.3 Analysis of Data Organization Paradigms
We compare various data organization paradigms within RAG, implementing Semantic Chunking (Kamradt, 2024) and Proposition (Chen et al., 2023) in addition to recursive chunking (Splitter, 2023) (chunk-based) and entire document (doc-based) approaches (details in Appendix A). Table 7 shows that our Thread paradigm outperforms the other paradigms across all metrics. Thread w/o. represents without LU selector. Although Semantic and Proposition methods use LLMs to merge semantically similar sentences, they fail to adequately address logical connections, resulting in poor performance. Additionally, our system processes the smallest token length of retrieval units yet achieves the highest performance, underscoring our approach’s efficiency and effectiveness.
Figure 4: Analysis of using different LLMs on WikiHow (GPT-4 v.s. LLaMA3-70B).
@@ 5.3 Analysis of Data Organization Paradigms
5.4 Document Formats in LU Extraction
As mentioned in §4.2, we test our LU extraction method in accommodating varying document structures. We generate different document formats (detailed in Appendix C.2), including structured markdown, hierarchical guidelines, tabular checklists, and narrative documents. Table 7 shows that our Thread paradigm effectively organizes these formats, consistently outperforming the chunk-based paradigm across all metrics. Our Thread improves Ele. Acc by up to 9.61%, Op. F1 by up to 8.43%, and Step SR by up to 8.51%. Especially, our Thread achieves the highest performance with structured documents, as this format helps construct a higher-quality knowledge base.
5.5 Different LLMs as backbone
We further conduct experiments on LLaMA3-70B, as shown in Figure 4. Although LLaMA3-70B is less powerful than GPT-4, it still demonstrates competitive performance with the help of Thread. Results indicate that LLaMA3 struggles with WikiHow questions applying chunk-based or doc-based paradigms. However, with the integration of Thread, LLaMA3 not only achieves much better performance in SR, but also narrows the gap with GPT-4. Specifically, LLaMA3 achieves SR and F1 scores of 60.82% and 72.84%, respectively, compared to GPT-4’s scores of 72.16% and 80.74%. This indicates that while a performance gap remains, particularly with the chunk-based and doc-based paradigms, our Thread considerably reduces this disparity between GPT-4 and LLaMA3. These findings highlight the value of our proposed paradigm in enhancing the performance of different LLMs, showcasing its generalizability, robustness, and efficiency in handling how-to questions.
@@ 6 Conclusion
In this paper, we address the overlooked category of handling how-to questions in QA systems by proposing Thread, a novel data organization paradigm that captures logical connections within documents. By introducing a new knowledge granularity called ’logic unit,’ Thread restructures documents into interconnected logic units that are compatible with RAG methods. Extensive experiments show that Thread significantly outperforms existing paradigms, improving performance while reducing the knowledge base size and minimizing the information needed for generation.
Appendix A Current Data Organization Paradigm
From Gao et al. (2023), current data organization paradigms can be categorized into phrases, sentences, propositions, chunks, and so on. In our paper, we choose chunks  [^9] We use the implementation by LangChain https://python.langchain.com/v0.2/ and propositions to compare with our proposed Thread  [^10] Note: For chunks, we retrieve the top-5 at each time, and for documents, we only retrieve the top-1..
Recursive Chunk. This chunking method splits the original documents using a list of separators, then reassembles them according to specified chunk sizes and overlap sizes. In our experiment, we use different chunk sizes for each dataset: 1000 for Mind2Web, 2000 for IcM, and 300 for WikiHow. The chunk overlap sizes also vary 50 for Mind2Web, 100 for IcM, and 30 for WikiHow.
Entire Document. This method sends the entire document directly into the model, constrained by the document’s length and structure.
Semantic Chunk. Kamradt (2024) proposes splitting chunks based on semantic similarity. The hypothesis is that semantically similar chunks should be grouped together. By comparing the semantic similarity between adjacent sentences, the method identifies “break points”. If the similarity in the embedding space exceeds a certain threshold, it marks the start of a new semantic chunk.
Agentic Chunk (Proposition). Chen et al. (2023) [^11] For proposition paradigm, we use agentic chunker since the input token of Flan-T5 is limited to 512, https://github.com/FullStackRetrieval-com/RetrievalTutorials. introduces the concept of the Proposition Paradigm, which involves extracting independent propositions from original documents. The Agentic Chunk method is based on this paradigm. It first splits the documents into paragraphs, then extracts propositions from each paragraph, and at last merges similar propositions into chunks.
In all our experiments, we set the temperature of LLMs to 0 and top_p to 1 for results reproduction.
Appendix B Example of Dataset
We list one example of each dataset in Table B, and an example of the construction process of Thread including the original document, reformulated document, and its corresponding logic unit in Table LABEL:tab:example_lu.
@@ Appendix B Example of Dataset
<html> ... </html> Based on the HTML webpage above, try to complete the following task Task: Book the lowest-priced and quickest flight for 5 adults and 1 child on May 20 from Mumbai to any airport near Washington. Previous actions: None What should be the next action? Please select from the following choices (If the correct action is not in the page above, please select A. None of the above’): A. None of the above B. <div id=0> <input radio triptype roundtrip true /> <label> <span> C. <label id=1> <span> Search flights one way </span> <span> One D. <a id=2> <h3> Celebrate World Wish Day </h3> <p> Support E. <h2 id=3> Help </h2> F. <a id=4> <img follow us on twitter - /> </a> C. Action: CLICK
WikiHow
@@
How to Investigate Service A-To-Service B Connection? ### Step 1: Check Pull Task Execution From the Cluster The direct impact of connection failure is pull task execution will not work. If Service A can continue to pull from Service B, then the incident can be dismissed as false alarm, the feature owner can investigate further to see why Echo fails. This can be visualized by pull task count over time in the last 8 hours in the following query: ‘‘‘kusto <Code Block> ‘‘‘ Disregard the last data point, if the data point is always above zero, then consider the alert as false alarm. If the chart sometimes drops to zero one hour ago and the number is low in general (for instance less than 20), it means the customer traffic in the cluster is low. In this case, observe for a longer period of time. If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted. Otherwise, continue to observe since Service A is pulling Service B just fine.
Table 8: Examples of each dataset. For Mind2Web, although the test set has fixed options for each step, there are different execution methods for the same task on each website, so it is essentially dynamic.
- Terminology: firstly, it should be the relationship or connection between terminology about the incident, if not, it can be the explanation or concept of the incident. Sometimes it should be extracted and summarized by yourself. - FAQ: frequently asked questions that help to understand the incident. - STEP: the processes to resolve the incident, and you should make sure its completeness. Usually, steps have causal inner connection, the former step will trigger the next step. - Appendix: the supplement of the incident that is not important or labeled by TSG, usually providing additional resources, data, links and so on.
1. You need to identify each second-level subsection, including third-level subsection if needed, analyze its content or purpose, and categorize it accordingly. For those belonging to Step, you should capture the inner connections, such as Causality or Temporal relations, and present them in the correct order.
2. Your returned formulated TSG should be in JSON format. Make sure that the keys originate from these categories: Terminology, FAQ, STEP ad Appendix. Each value should be a list of dictionaries. The keys for them are "prerequisite", "header", "body", and "linker". All values within the lists need to align with the original context, with truthful meaning and necessary **code block**.
3. Importantly, the "linker" is used to imply the dual role of providing the action’s result and connecting to the next step using the "if-then" sentence format. You should formulate each step’s linker to be "If any results are obtained by executing the corresponding action in the previous step, then **the true intent of the following step** provided here". Implicit linkers like "proceed to the next step." or "then the intent of the following step should be taken into consideration." should be avoided.
4. For each "if" condition at every step in the STEP, it is necessary to add a special token behind the "then" condition within the "linker". The options for these tokens are "[CONTINUE]", "[CROSS]", and "[MITIGATE]". - The token "[CONTINUE]" indicates that the actions corresponding to this "if" condition are part of the continuum within the same TSG’s STEPs. - The token "[CROSS]" signifies that the subsequent actions require a transition to a different set of steps that are external to the current TSG’s STEPs. - The token "[MITIGATE]" implies that the actions following the "if" condition convey that the incident is mitigated, or necessitate communication with on-call engineers or teams.
The use of this special token is instrumental in verifying the completeness and structural integrity of the STEP section.
<TWO EXAMPLES HERE>
[User]
Here is the <TSG> you need to formulate: {TSG}
Instruction that formulates the original unstructured troubleshooting guide into structured one.
<ONE EXAMPLE HERE>
Your response should be in the JSON format as below:
"#CODE_TEMPLATE#": where you replace the parameters in <CODE> with placeholders,
"#DEFAULT_PARAMETERS#": where you keep the parameters in <CODE> as default values.
[User]
Here is the <CODE> you need to extract: {CODE}
Instruction that extracts code template and default parameters from the source code.
Each element in <LU_LIST> is in json format and contains the following fields:
"#type#": "the type of the element, select from the following types: Terminology, FAQ, Step, and Appendix.",
"#meta data#": "the description of the troubleshooting guide.",
"#prerequisite#": "The prerequisite of this step, before taking the current step, the prerequisites should be finished.",
"#header#": "The information describes the intent of the <INFO>.",
"#body#": "The action is the content which troubleshoots the incident or explain the #header#. the action may contain code blocks in markdown format, and parameters are replaced with placeholders",
"#linker#": "the expected output after taking the #action#. It is defined in the following format in markdown: -If **condition**, then **should_do**. It can contain multiple if-then cases.",
"#default_parameters#": "the default parameters that could fill in placeholders in code blocks in #body#."
- The elements in <LU_LIST> contain possible information that can answer the user’s query in <QUERY>. However, they may not be all relevant to the query or useful to answer the user’s query. You should select the most relevant element from the <LU_LIST> based on the user’s query in <QUERY>.
- In particular, you should focus on the following fields in the element: #header#, #body#. Most importantly, the <QUERY> needs to match with the #intent# and the #body# has to provide actions to reach the goal of the <QUERY>, please ignore the #linker# and do not map the <QUERY> with #linker#.
- As you choosing from <LU_LIST>, you need to check if all the #prerequisite# are met in previous history. If the #prerequisite# is not finished, then it should not be chosen.
- Try to select only one element from <LU_LIST>. If it is not possible to select only one element, you can select multiple elements from <LU_LIST>:
"INDEX": the index of the element in <LU_LIST>.
"INTENT": the #header# of the element, the index starts from 0.
"EXPLANATION": justify why you select this node.
- If there is no element in <LU_LIST> that can answer the user’s query in <QUERY>, you should try to select the most relevant element to the user’s query considering that the user might use the wrong terminology:
"INDEX": the index of the element in <LU_LIST>.
"INTENT": the #header# of the element, the index starts from 0.
"REPHRASED_QUERY": the rephrased query that you think the user is asking about.
"EXPLANATION": justify why you select this node.
- Unless you are confident that there is no element in <LU_LIST> that is even close to the user’s query:
"NO_INFO_EXPLANATION": where you give your explanation.
- Your answer should be in the JSON format in a list after <RESPONSE>.
[User]
<LU_LIST>: {LU_LIST}
<QUERY>: {QUERY}
<CHAT_HISTORY>: {CHAT_HISTORY}
Instruction that selects the most relevant logic unit based on user query and chat history.
C.2 Mind2web
We show the details about the document generation instruction we use, the different formats of documents, and the examples we generate.
The data includes a step-by-step execution process, each step encompassing HTML code, Tasks, Previous Actions, and the Element and Action of this step. Note that the Element comes from the HTML code, and if the correct action is not present on the current page, the Element is None, and you should retrieve it from next step’s Previous Actions.
Now your task is to write a comprehensive and adaptable reference document that outlines the general process for completing tasks like the given task. This document should serve as a guide for others to perform similar tasks on the same website in the future. So it should not be limited but can use this data to be the example, and should be general enough.
Please return the complete reference document that adheres to these guidelines.
[User]
The format of the documents should be as follows: {FORMAT}
The given execution process is as follows: {EXAMPLE}
Instruction that generates specific format of document for Mind2web dataset..
We follow Wang et al. (2024) to brainstorm diverse formats of documents for Mind2web dataset used for retrieval, and the results are listed in Table 10.
Format Description
Structured
Markdown
- The document must be structured into sections in markdown format. - It should include a task overview, introduction, process steps, and conclusion. - Each step in the process includes detailed explanations for Intent, Prerequisite, HTML Code Reference, Action, Reason, and Result. - The Prerequisite is to specify any conditions or prior actions that must be met or completed before proceeding with the current step in the process. - Ensure that each step is explicitly connected to the next one, and the result is written in the "if-then" schema where the "Intent" of this step is completed, and the outcome "then" is the next step’s Intent. - The HTML Code Reference gives hints of the Action like some ’<button>’, ’<span>’, or other elements or attributes. You need to use the given task as an example. - The Action comes from "Click", "Type", "Hover", "Press Enter".
Hierarchical Guideline - A structured text document with numbered steps for each task. - Each step includes a title, description, the HTML code involved, and the action to be taken. - Previous actions are referenced where necessary, with hyperlinks to the relevant steps. - Appendices for HTML code references, glossary of terms, and FAQs.
Tabular
Checklist
- A printable checklist with each task and subtask, including checkboxes for completion. - Each checklist item includes a code snippet and the action required. - A troubleshooting section that lists common problems and their solutions. - Tips for what to do when the expected element or action is not available. - References to more detailed instructions or external resources for complex tasks.
Narrative
Document
An entire description of the execution process without special structures.
Table 10: The description of different formats of documents on Mind2web.
C.3 Wikihow
I will give you the answer generated by the system in <Generation> and the ground truth answer in <Ground Truth> respectively. Your evaluation will contain five sub-evaluation tasks:
1. Both two answers contain a list of steps. Your task is to extract action items from the provided steps in both answers. The action item is defined as a combination of action and element. Compare the action items to identify similarities. Output the similar action items. Count the count of similar action items.
- Your answer should contain the extracted two action item sets (in the format as a list of strings). - Your answer should contain a set of similar action items (in the format of a list of strings). Similar action items are those sharing similar intent or achieving similar goals. Each similar action pair in the list should be in the format of "similar action item from action item set1 / similar action item from action item set2" - Your answer should contain the count of similar action items.
2. Can <Generation> completely solve the user’s question? - Your answer should be "Yes" or "No". - Your answer should contain the reason(s) for your choice. You should not focus on the length of the answer or the details of the answer, but you should focus on whether the steps could solve the user’s question and the quality of the steps compared with the ground truth.
Your output should be in the following format in JSON:
"Subtask1": {
"Action items in Generation": ["action item 1", "action item 2", ...],
"Action items in Ground Truth: ["action item 1", "action item 2", ...],
"Similar action items": ["similar action item 1", "similar action item 2", ...],
"Count of similar action items": 2
"Subtask2": {
"Choice": "Yes" or "No",
"Reason": "reason for your choice"
[User]
Here is the user’s question <Question>: {Question} The answer from system <Generation> is: {Generation} The ground truth answer <Ground Truth> is: {Ground Truth}
Instruction that evaluates the generated answer compared with ground truth for Wikihow.
Appendix D Details about RAG System
We take the scenario of Mind2web to show the instructions we use in our RAG-based QA system.
Your response should be in the format of "Answer: C. Action: SELECT Value: Pickup". The answer is A, B, C…, the Action comes from [CLICK, TYPE, SELECT] and the Value is not always needed.
[User]
<DOC_CONTEXT>: {DOC_CONTEXT} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
Instruction that is used for the baselines of RAG system on Mind2web.
Your response should be in the format of JSON:
"CHOICE": the choice you make from A B C ...,
"ACTION": the corresponding action choosing from [’CLICK’, TYPE’, SELECT’],
"VALUE": the corresponding value if needed,
"INTENT": the intent of the next step, which should be retrieved and judged from the "if" conditions in #output# from <LU> according to the current step and actions and choose the corresponding "then" outcome, do not guess it based on current Task in <CURRENT_STEP> by yourself unless the <LU> is irrelevant to <CURRENT_STEP>,
[User]
<LOGIC_UNIT>: {LOGIC_UNIT} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
Instruction that is used for the RAG system utilizing Thread Paradigm on Mind2Web.
Appendix E Limitations
This work focuses on evaluating the effectiveness of Thread by designing how-to questions in three specific scenarios, covering both linear and dynamic how-to questions. However, there are several limitations that point to future directions. First, as our logic-based knowledge base can coexist with the original chunk-based knowledge base, we do not extend our method to factoid questions, such as multi-hop or long-form questions. Second, the experiments primarily use ChatGPT and LLaMA as the backbones. Future work should include evaluations with other LLMs and alternative retrievers such as Contriever to further validate the effectiveness of Thread. Lastly, while extracting logic units involves an initial cost in terms of LLM usage, this is a one-time process. Once the knowledge base is constructed, it provides significant advantages for industrial applications, particularly in terms of subsequent updates and maintenance.

### v3

@@ 3.2 Baselines and Metrics
Previous work on the Mind2Web dataset, such as MINDACT (Deng et al., 2023a), has not approached it as how-to questions, instead using In-Context Learning (ICL) or Supervised Learning (SL). For fair comparison with our LLM endpoints, [^6] We use GPT-3.5 and GPT-4 (version 1106-preview). we re-implement MINDACT using the same demonstrations and include chat history as context. In this work, we treat Mind2Web as dynamic how-to questions and address them with our RAG system (Appendix LABEL:appendix_rag shows how we adapt Mind2Web into how-to questions.). We compare Thread against document-based and chunk-based data organization paradigms, also used as baselines for the WikiHow and IcM datasets. Details on evaluation metrics, experimental setup (prompts, etc.), and data organization paradigms are provided in Appendix LABEL:appendix_metrics, LABEL:appendix_exp, and LABEL:appendix_paradigms, respectively.
@@
Table 2: Experimental results on Mind2Web. ‘*’ represents taking results from the original paper. Deng et al. (2023a) formulate web navigation as a series of multiple-choice steps, where each step requires selecting the correct HTML code and actions. SR refers to the overall Success Rate.
3.3 Main Results
Web Navigation
Table 2 presents the overall performance on Mind2Web, comparing our method with baselines and both doc-based and chunk-based RAG methods. We find that incorporating informative documents, regardless of the data organization paradigm, significantly improves performance. RAG methods outperform both ICL and SL approaches, with the doc-based RAG method achieving results comparable to MINDACT’s best performance. Notably, Thread surpasses all RAG baselines, improving Ele. Acc by 4.06%, Step SR by 3.49%, and SR by 3.57%. MINDACT-SL achieves the highest Op. F1 due to label imbalance [^7] Predicting all operations as “CLICK” results in an Op. F1 of 79.90%., biasing the model toward frequent operations.
@@ 3.3 Main Results
Paradigm SR Step SR P.F. Step SR HI Turns
40.51 60.90 60.90 30.10 3.14
Chunk 28.95 53.16 43.05 46.84 6.84
43.86 63.90 63.90 36.09 2.98
Doc 31.58 57.89 42.11 42.11 6.53
77.19 88.72 84.21 11.28 2.56
Thread 52.63 84.21 68.95 15.79 5.74
Table 3: Experimental results on Incident Mitigation. Incidents are divided into simple and hard groups. P.F. Step SR: the perctenage of successful steps before the first failure. HI: the steps requiring human intervention.
Incident Mitigation
Table 3 highlights the advantages of Thread in handling complex dynamic how-to questions. Both chunk-based and doc-based RAG methods struggle with incident mitigation, as their low SR and P.F. Step SR scores indicate difficulty in connecting subsequent steps based on the current state. In contrast, Thread achieves the highest performance across all metrics, notably improving SR from 21.02% to 33.33%. More importantly, its superior P.F. Step SR demonstrates its ability to dynamically link steps based on user feedback, reducing human intervention. As a result, Thread not only achieves the highest Step SR but also minimizes interaction turns, effectively mitigating both simple and complex incidents.
Wikipedia Instructions
Table 4 compares different data organization paradigms on WikiHow under two settings. For single-turn, where the RAG system generates the entire plan in one step, Thread outperforms the doc-based by 5.15%, providing more concise and effective information. In contrast, the chunk-based achieves a significantly lower SR of 19.59%, highlighting the challenge of retrieving relevant chunks without maintaining the document’s logical flow. For multi-turn, iterative retrieval improves performance over single-turn, demonstrating the benefits of a step-by-step approach for how-to questions. The SR increases from 58.76% to 68.04% for the doc-based and from 63.91% to 72.16% for Thread. As in single-turn case, the chunk-based method disrupts internal logic, resulting in a low SR of 20.62%. Overall, Thread achieves the highest SR of 72.16%, emphasizing its effectiveness in preserving and modeling step dependencies. We also extend our experiments to LLaMA3-70B to show the generalibity of Thread in Appendix LABEL:sec:backbone.
@@
Table 4: Experimental results on WikiHow with different interaction manners and paradigms.
4 Analysis
@@ 4.1 Ablation on RAG System Settings
This section presents an ablation study on key components of our RAG system, using the Mind2Web dataset. While retriever and generator variants have been explored in prior work (Gao et al., 2023), we utilize text-embedding-ada-002 (OpenAI, 2022) for retrieval and GPT-4 as the generator.
@@
Multi-turn Interaction. As shown in Table 4, the multi-turn setting outperforms the single-turn setting in answering how-to questions. We adopt the multi-turn setting for all scenarios.
Chat History. Chat history enables the system to reference previous actions and results, improving decision-making. Table 5 shows that excluding chat history results in performance degradation. Therefore, we include chat history for all RAG-based methods in our experiments.
Retrieval Units Selector. As described in §2.4, the Selector identifies the most relevant retrieval units from the top-K retrieved units. Table 5 illustrates the effect of the retrieval unit selector. Without the selector, performance drops by 4.29% in Ele. Acc and 3.38% in Op. F1. The selector improves all metrics when applied to Thread. Unlike LU selection, which filters irrelevant LUs based on prerequisites, chunk selection ignores inter-chunk connections and may exclude relevant chunks. Thus, we activate the selector only for Thread.
4.2 Comparison with Different Paradigms
We compare several RAG data organization paradigms (details in Appendix LABEL:appendix_paradigms), including Semantic Chunking (Kamradt, 2024), Proposition (Chen et al., 2023), recursive chunking (Splitter, 2023) (chunk-based), entire document (doc-based), and GraphRAG (Edge et al., 2024). As shown in Table 6, Thread outperforms all baselines across metrics. While Semantic and Proposition merge semantically similar sentences using LLMs, they fail to capture logical relationships. And GraphRAG underperforms on how-to questions due to its focus on entity-level relations over document logic. In contrast, Thread retrieves smaller, logic-driven units, yielding higher efficiency and accuracy. Appendix LABEL:case_study illustrates the difference when handling the question in Table LABEL:tab:example_lu between GraphRAG and Thread, with GraphRAG producing flawed outputs. Appendix LABEL:appendix_result analyzes Thread ’s cost and scalability, confirming its high performance at acceptable costs and potential for large-scale datasets.
@@ 4.2 Comparison with Different Paradigms
4.3 Superiority over Different Doc Formats
As real-world documents vary in format, we test our LU extraction method on diverse structures, including structured markdown, hierarchical guidelines, tabular checklists, and narrative documents (details in Appendix LABEL:appendix_mind2web). Table 7 demonstrates that Thread consistently outperforms the chunk-based paradigm across all metrics, improving Ele. Acc by up to 9.61%, Op. F1 by up to 8.43%, and Step SR by up to 8.51%. These results underscore Thread’s ability to effectively handle a variety of real-world document resources. Notably, Thread achieves the highest performance with structured documents, which facilitate the creation of a higher-quality knowledge base.
@@ 4.3 Superiority over Different Doc Formats
Figure 3: Experimental results on the generalization of Thread to 2WikiMultiHopQA.
4.4 Generalization to Open Domain Questions
To explore whether Thread can be extended to open-domain questions, we evaluate Thread on multi-hop questions, 2WikiMultiHopQA (Ho et al., 2020). Specifically, we sample 200 questions of two types, ‘inference’ and ‘compositional’ type, like “Who is the maternal grandfather of Abraham Lincoln?” and “Who is the founder of the company that distributed La La Land film?” In our experiment, we use the relations between entities as ‘Linker’. The findings from Figure 3 demonstrate that THRED largely surpasses the chunk-based paradigm on multi-hop questions, validating the generalization of Thread on diverse questions.
5 Related Work
Data Organization Paradigm in RAG
The data organization process is a critical pre-stage of RAG methods where documents are segmented following certain data organization paradigms. The most common data organization paradigm is splitting documents into retrieval units (Gao et al., 2023). These retrieval units vary in granularity such as phrases, sentences, propositions (Chen et al., 2023), chunks (Kamradt, 2024), etc. Coarser-grained units contain more information but introduce redundant noise, while finer-grained units have lower semantic integrity and often require retrieving more units to gather comprehensive information. However, the chunk-based data organization paradigm ignores the logical and relational connections between chunks, potentially disrupting the inherent logic flow in documents. Another paradigm constructs documents into knowledge graphs (KG), where retrieval units include entities, triplets, etc. (Gaur et al., 2022; Sen et al., 2023; He et al., 2024; Wang et al., 2024). While these approaches emphasize semantic and lexical similarities between retrieval units, their effectiveness on how-to questions remains limited. This is largely due to the inherent difficulty in capturing the logical structure required by such questions. In particular, current methods such as GraphRAGEdge et al. (2024), primarily model relationships between different chunks, which fails to account for the more complex, logic-driven connections between chunks that go beyond entity-level similarity.
Information Retrieved by RAG
@@ 6 Conclusion
In this paper, we address the overlooked category of handling how-to questions in QA systems by proposing Thread, a novel data organization paradigm that captures logical connections within documents. By introducing a new knowledge granularity called ‘logic unit’, Thread restructures documents into interconnected logic units that are compatible with RAG methods. Extensive experiments show that Thread significantly outperforms existing paradigms, improving performance while reducing the knowledge base size and minimizing the information needed for generation.
Limitations
This work focuses on evaluating the effectiveness of Thread by designing how-to questions in three specific scenarios, covering both linear and dynamic how-to questions. However, several limitations point to future directions. First, as our logic-based knowledge base can coexist with the original chunk-based knowledge base, we only extend our method to multi-hop questions, more open-domain tasks should be considered. Next, while extracting logic units involves an initial cost in terms of LLM usage, this is a one-time process. Once the knowledge base is constructed, it provides significant advantages for industrial applications, particularly in terms of subsequent updates and maintenance.
Acknowledgement
We thank all reviewers for their great efforts. This work is supported by the National Science Foundation of China under Grant No.61876004, which covered personal research efforts, conference registration fees of K.K. An, S.Z. Si and B.B. Chang.
Appendix A Scenarios and Datasets
A.1 Documents under Each Scenario
As the Mind2Web dataset lacks relevant documents, we create retrieval documents tailored to the RAG system. Assuming that each website has help docs applicable across different tasks, we use the "Cross-Task" test set, selecting examples from the training set to craft informative documents for each website. And we follow Wang et al. (2023) to brainstorm different formats of documents. More details about collecting documents are shown in Appendix LABEL:appendix_mind2web.
For the IcM dataset, we collect 56 TSGs from an enterprise-level engineering team responsible for a large-scale cloud platform. The selected incidents in the IcM dataset can be resolved using the knowledge provided in these TSGs.
A.2 Example of Each Dataset
We list one example of each dataset in Table LABEL:tab:example_dataset, and we distinguish how-to questions into ‘linear’ and ‘dynamic’ types, where linear how-to questions often involve a fixed sequence of steps that do not require feedback or decision points based on intermediate outcomes.
@@ A.2 Example of Each Dataset
<html> ... </html> Based on the HTML webpage above, try to complete the following task Task: Book the lowest-priced and quickest flight for 5 adults and 1 child on May 20 from Mumbai to any airport near Washington. Previous actions: None What should be the next action? Please select from the following choices (If the correct action is not in the page above, please select A. None of the above’): A. None of the above B. <div id=0> <input radio triptype roundtrip true /> <label> <span> C. <label id=1> <span> Search flights one way </span> <span> One D. <a id=2> <h3> Celebrate World Wish Day </h3> <p> Support E. <h2 id=3> Help </h2> F. <a id=4> <img follow us on twitter - /> </a> C. Action: CLICK WikiHow